A very long wire generates a magnetic field of 0.0020 x 104 T at a distance of 1.0 mm. What is the magnitude of the current? (1mm- 10 m. A) 1.0 mA B) 20 mA C) 4,000 mA D) 3,100 mA 1. A conductor is formed into a square of side 0.5m and placed in the horizontal y-plane. A magnetic ficld of IT is oriented making 30.0P with the plane of the coil. What is the magnetic flux through the conductor? A) 0.12T-m2 B)022T-m2 C)025T-m2 D) 20T-to2 2. s A device with a resistance of 200.o k is connected to a 10.0 V battery. How much poerdoesthedence use? (lke 10kilili-10.) A)050W B) 200000k C) 0050 D) 2000W
Answer #1

1.

Given data

B-0.002 × 10-4T

r=1mm=10^{-3}m

We know

B=\frac{\mu _{o}}{2\pi }\frac{I}{r}

I=\frac{2\pi Br}{\mu _{o}}

I=\frac{2\pi \times 0.002\times 10^{-4}\times 10^{-3}}{4\pi \times 10^{-7}}A

I=1mA

Hence A) is the right answer.

2.

Given data

B=1T

A=(0.5\times 0.5)m^{2}=0.25m^{2}

Angle to the normal to that plane can be given by

\theta =60^{0}

Electric flux through that plane can be given by

\phi =BAcos\theta

\phi =(1\times 0.5\times 0.5\times cos60^{0})T.m^{2}

\phi =0.12T.m^{2}

Hence A) is the right answer.

3.

Given data

R=200k\Omega =200\times 10^{3}k\Omega

V=10V

Power used by the device can be given by

P=\frac{V^{2}}{R}

P=\frac{10^{2}}{200\times 10^{3}}W

P=0.05mW

Hence C) is the right answer.